2. Download the annual real GDP and GDP data of the United States 1950-2018 from FRED. For the real GDP, the data online is cShort Answer Questions (5 Points Each) 1. Based on BEA records, US GDP data from 2012 and first quarter of 2013 is given belo1. A small country has a population of 10,000. 8,000 are the civilian non-institutional population. 2,000 are not in labor fo4. If Jim used to work at a typewriter factory, and the popularity of personal computers results in the factory being shut doConsider the following data for an imaginary economy. Price Quantity Produced 10 Goods Shoes Shirts Cotton 20 40 1 7. Suppose9. Pick the appropriate option. Statement (i) The CPI market basket contains all goods and services produced in an economy. (15. Suppose your nominal income increases from $20,000 in year 1 to $25,000 in year 2. If the CPI was 100 in year 1, how much
Answer #1

(First question - depends on your Buck ID which you have not provided. Question cannot be answered)

(Short Answer Question 1)

For QI-2013:

(1) C ($B) = Consumption of goods + Consumption of services = 3,873 + 7,500 = 11,373

(2) I ($B) = Fixed investment + Change in inventories = 2,103 + 48 = 2,151

(3) G ($B) = Federal expenditure + State and local expenditure = 1,586 + 1,846 = 3,432

(4) (X - M) ($B) = Exports - Imports = 2,204 - 2,747 = - 543

(5) GDP ($B) = C + I + G + (X - M) = 11,373 + 2,151 + 3,432 - 543 = 16,413

NOTE: As HOMEWORKLIB answering Policy, 1st question is answered.